The food compartment of a refrigerator, is maintained at 4°C by
removing heat from it at a rate of 360 kJ/min. If the required power input
to the refrigerator is 2 kW, determine (a) the coefficient of performance of
the refrigerator and (b) the rate of heat rejection to the room that houses
the refrigerator.
1
Expert's answer
2018-11-15T09:42:09-0500

Let us convert the cooling load (Q) of the refrigerator to kW:

Q=360 kJ/min=6 kW.

The coefficient of performance (COP) of the refrigerator is given by the ratio of the cooling load and input power (W). Thus:

COP=Q/W=6/2=3.

The rate of heat rejection (H) to the room is equal to the sum of the refrigerator cooling load and input power:

H=Q+W=6+2=8 kW.